The sum of two numbers is 76. The larger number is 14 more than the smaller number. What are the two numbers?

Found 2 solutions by Fombitz, mathmate:
Answer by Fombitz(32388)

Duplicate Question
See Answer 700483

Answer by mathmate(429)

Question:
The sum of two numbers is 76. The larger number is 14 more than the smaller number. What are the two numbers?

Solution:
Sum and difference problem, many of us can do this in the head.
Larger number = (sum+difference)/2=(76+14)/2=45
Smaller number = (sum-difference)/2=(76-14)/2=31

Can also solve by system of equations:
L=larger number
S=smaller number
L+S=76 ...........(1)
L-S=14 ...........(2)
Add (1) & (2)
2L=76+14=90
L=90/2=45
Subtract (2) from (1)
2S=76-14=62
S=62/2=31